\newpage
\begin{thebibliography}{99}

\bibitem{Bax} R.J.Baxter: Exactly solved models in statistical
mechanics. Academic Press, 1982
\bibitem{CG} L.Alvarez-Gaum\'e, C.G\"mez and G.Sierra,
Nucl. Phys. B319 (1989) 155, Phys. Lett. B220 (1989) 142, Nucl.
Phys. B330 (1990) 347

G.Moore and N.Yu.Reshetikhin, Nucl. Phys. B328 (1989) 557

C.G\'omez and G.Sierra, Nucl. Phys. B352 (1991) 791
\bibitem{FR} I.B.Frenkel and N.Yu.Reshetikhin, Commun. Math.
Phys. 146 (1992) 1
\bibitem{SR} L.Rozansky and H.Saleur, Nucl. Phys. B376 (1992) 461

L.H.Kauffman and H.Saleur, Commun. Math. Phys. 141 (1991) 293

\bibitem{ACG} A.Berkovich, C.G"mez and G.Sierra,``Spin-Anisotropy
Commensurable Chains: Quantum Group Symmetries and N=2 SUSY",
preprint IMAFF-93-1(1993)

\bibitem{B} D.Bernard and A.LeClair, Phys. Lett. B247 (1990) 309
\bibitem{FI} P.Fendley and K.Intriligator, Nucl. Phys. B380 (1992)
265
\bibitem{FW} C.Fan and F.Y.Wu, Phys. Rev. B2 (1970) 723
\bibitem{F} B.U.Felderhof, Physica 66 (1973) 279
\bibitem{BS} V.V.Bazhanov and Yu.G.Stroganov, Teor. Mat. Fiz. 62
(1985) 337
\bibitem{LSM} E.H.Lieb, T.D.Schultz and D.C.Mattis, Ann.
Phys.(NY) 16 (1961) 407
\bibitem{R} H.Hinrichsen and V.Rittenberg, Phys. Lett. B275
(1992) 350
\bibitem{S} E.K.Sklyanin, Func. Anal. Appl. 16 (1982) 263, Func. Anal. Appl.
17 (1983) 273
\bibitem{Ch} I.V.Cherednik, Yad. Fiz. 36 (1982) 549, Fun. Anal.
Appl. 19 (1985) 77

A.V.Odeski and B.L.Feigin, Func. Anal. Appl. 23 (1989) 207
\bibitem{Z} A.B.Zamolodchikov, Commun. Math. Phys. 69 (1979) 165
\bibitem{ZDV} A.B.Zamolodchikov, Sov. J. of Nucl. Phys. 46
(1987) 1090

C.Destri and H.J.de Vega, Mod. Phys. Lett. A4 (1989) 2595
\bibitem{T} M.Toda, Phys. Rep. 18 (1975) 1

\end{thebibliography}

==================

\documentstyle[12pt]{article}

\begin{document}

\begin{titlepage}
\title{The Hidden Quantum Group of the 8--vertex Free Fermion
Model: q--Clifford Algebras}

\author{R. Cuerno, C. G\'omez, E. L\'opez--Manzanares and G. Sierra \\
{\it Instituto de Matem\'aticas y F\'{\i}sica Fundamental, CSIC} \\
{\it Serrano 123, E--28006 Madrid, SPAIN}}

\date{}

\maketitle

\begin{abstract}
We prove in this paper that the elliptic $R$--matrix of the
eight vertex free fermion model is the intertwiner
$R$--matrix of a quantum deformed Clifford--Hopf algebra. This
algebra is constructed by affinization of a quantum Hopf
deformation of the Clifford algebra.
\end{abstract}

\vskip-12.0cm
\rightline{{\bf IMAFF-2/93}}
\rightline{{\bf February 1993}}
\vskip2cm

\end{titlepage}

\section{Introduction}

The realm of two dimensional integrable models contains two
important families associated to the six vertex and eight vertex
solutions to the Yang-Baxter equation \cite{Bax}.

Whereas the family of six vertex solutions (six vertex model and
their higher spin descendants) are $R-$matrix intertwiners for
different finite dimensional irreducible representations of
$U_q(\widehat{Sl(2)})$, the elliptic eight vertex solutions do not
admit, for the time being, the interpretation as quantum group
intertwiners. To find such a quantum group interpretation of the
eight vertex model would provide, for instance, a natural way to
extend the known hidden quantum group structure of conformal
field theories \cite{CG} to $q-$conformal field theories defined by the
$q-$deformed Knizhnik-Zamolodchikov equation \cite{FR}.

A special class of solutions to the vertex Yang-Baxter (YB) equation
are the ones satisfying the so called free fermion condition:
\begin{equation}
R_{00}^{00}(u)R_{11}^{11}(u)+R_{01}^{10}(u)R_{10}^{01}(u)=
R_{00}^{11}(u)R_{11}^{00}(u)+R_{01}^{01}(u)R_{10}^{10}(u)
\label{2}
\end{equation}

In the six vertex case, $R_{00}^{11}(u)\!=\!R_{11}^{00}(u)\!=\!0$
the solutions to YB satisfy (\ref{2}) and are given by the $R-$matrix
intertwiners of the Hopf subalgebra $U_{\hat q}(\widehat{gl(1,1)})$
\cite{SR}. These intertwiners can be mapped into the ones
of $U_q(\widehat{sl(2)})$ ($q^4 \!=\!1$) for non classical nilpotent
irreducible representations \cite{ACG} with $\hat{q} \!=\! \lambda$
and $\lambda^2$ the eigenvalue of the casimir $K^2$. The
physical interest of the free fermion six vertex solutions is
their close connection with $N\!=\!2$ integrable models. In fact we
can define using the generators of $U_q(\widehat{sl(2)})$ for $q^4 \!=\!1$
a $N\!=\!2$ supersymmetric algebra \cite{B}, and in this case
the free fermion condition
(\ref{2}) reflects the $N\!=\!2$ invariance of the $R-$matrix.
Moreover  the $N\!=\!2$ piece of the solitonic $S-$matrix for the $N\!=\!2$
Ginzburg-Landau superpotential $W\!=\!X^{N+1}/(N+1)\!-\! \beta X$
\cite{FI} can be shown to be given by the intertwiners of
$U_{\hat{q}}(\widehat{gl(1,1)})$ with $\hat{q}^N \!=\!1$.

In the eight vertex case, $R_{00}^{11}(u) \! \neq \!0$
$R_{11}^{00}(u) \! \neq \!0$, solutions to YB satisfying (\ref{2}) have
been known for a long time \cite{FW}. The most general solution
corresponding to imposing non-zero field \cite{F}, \cite{BS}
depends on three spectral parameters $u$, $\psi_1$, $\psi_2$,
and is given by:
\begin{eqnarray}
a&\equiv &R_{00}^{00}=1-e(u)e(\psi_1)e(\psi_2) \nonumber \\
\tilde{a}&\equiv &R_{11}^{11}=e(u)-e(\psi_1)e(\psi_2) \nonumber \\
b&\equiv &R_{01}^{10}(u)=e(\psi_1)-e(u)e(\psi_2) \label{3} \\
\tilde{b}&\equiv &R_{10}^{01}=e(\psi_2)-e(u)e(\psi_1) \nonumber \\
c&\equiv &R_{01}^{01}=R_{10}^{10}=(e(\psi_1)sn(\psi_1))^{1/2}
(e(\psi_2)sn(\psi_2))^{1/2}(1-e(u))/sn(u/2) \nonumber \\
d&\equiv &R_{00}^{11}=R_{11}^{00}=-ik(e(\psi_1)sn(\psi_1))^{1/2}
(e(\psi_2)sn(\psi_2))^{1/2}(1+e(u))sn(u/2) \nonumber
\end{eqnarray}

\noindent
with $e(u)$ the elliptic exponential:
\begin{equation}
e(u)=cn(u)+isn(u)
\label{4}
\end{equation}

\noindent
and $k$ the elliptic modulus. The Yang--Baxter equation
satisfied by this $R$ matrix is \cite{BS}:
\begin{eqnarray}
& & ({\bf 1} \otimes R(u;\psi_1,\psi_2))(R(u+v;\psi_1,\psi_3)
\otimes {\bf 1})({\bf 1} \otimes R(v;\psi_2,\psi_3)) = \nonumber \\
& & (R(v;\psi_2,\psi_3) \otimes {\bf 1})({\bf 1} \otimes
R(u+v;\psi_1,\psi_3))(R(u;\psi_1,\psi_2) \otimes {\bf 1})
\label{1}
\end{eqnarray}

\noindent
The simplest way
to catch the physical meaning of solution (\ref{3}) is to define
the corresponding spin chain hamiltonian:
\begin{equation}
\left. H= \sum_{j=1}^{N}i \frac{\partial}{\partial u}R_{j,j+1}(u;\psi,\psi)
\right|_{u=0}
\label{5}
\end{equation}

\noindent
which is the well known $XY-$ model in an external magnetic
field \cite{LSM}:
\begin{equation}
H= \sum_{j=1}^{N} [ (1+ \Gamma) \sigma_j^x \sigma_{j+1}^x +
(1- \Gamma) \sigma_j^y \sigma_{j+1}^y + h(\sigma_j^z +
\sigma_{j+1}^z) ]
\label{6}
\end{equation}

\noindent
where:
\begin{eqnarray}
\Gamma & = & \frac{2cd}{ab+ \tilde{a} \tilde{b}} = ksn(\psi)
\label{7} \\
h & = & \frac{a^2+b^2- \tilde{a}^2- \tilde{b}^2}{2(ab+\tilde{a}
\tilde{b})}=cn(\psi) \nonumber
\end{eqnarray}

In this letter and as a preliminary step of the long term
process of finding the quantum group symmetry of the eight vertex
model, we will define a fully fledged Hopf algebra such that
its $R-$intertwiners coincide with the elliptic free fermionic
eight vertex solution (\ref{3}).

\section{The quantum Clifford algebra}

A Clifford algebra $C(\eta)$ associated to a cuadratic form
$\eta$ is the associative algebra generated by the elements $\{
\Gamma_{\mu} \}_{\mu =0}^D$, which satisfy:
\begin{equation}
\{ \Gamma_{\mu}, \Gamma_{\nu} \} = 2 \eta_{\mu \nu} {\bf 1}
\;\;\; \mu = 1,\ldots,D
\label{8}
\end{equation}

Associated to $C(\eta)$ we define the Clifford--Hopf algebra
CH(D) as the associative algebra generated by $\Gamma_{\mu}
\;\;\;(\mu=1,\ldots,D),\;\; \Gamma_{D+1}$ and the central elements
$E_{\mu} \;\;\; (\mu=1,\ldots,D)$ satisfying the following relations:
\begin{eqnarray}
& & \Gamma_{\mu}^2 = E_{\mu} \;\; , \;\; \Gamma_{D+1}^2 = {\bf 1}
\nonumber \\
& & \{ \Gamma_{\mu}, \Gamma_{\nu} \} =0, \;\; \mu \neq \nu
\nonumber \\
& & \{\Gamma_{\mu}, \Gamma_{D+1} \} =0 \label{9} \\
& & [ E_{\mu}, \Gamma_{\nu} ] = [ E_{\mu}, \Gamma_{D+1} ] =
[ E_{\mu}, E_{\nu} ] = 0 \;\; \forall \mu, \nu \nonumber
\end{eqnarray}

\noindent
The algebra CH(D) is a Hopf algebra with the following
comultiplication $\Delta$, antipode $S$ and counit $\epsilon$:
\begin{equation}
\begin{array}{lll}
\Delta (E_{\mu}) = E_{\mu} \otimes {\bf 1} + {\bf 1} \otimes
E_{\mu}, & S(E_{\mu}) = -E_{\mu}, &\epsilon(E_{\mu}) = 0 \\
\Delta (\Gamma_{\mu}) = \Gamma_{\mu} \otimes {\bf 1} + \Gamma_{D+1}
 \otimes \Gamma_{\mu}, & S(\Gamma_{\mu}) = \Gamma_{\mu}
\Gamma_{D+1}, & \epsilon(\Gamma_{\mu}) = 0 \\
\Delta(\Gamma_{D+1}) = \Gamma_{D+1} \otimes \Gamma_{D+1}, &
S(\Gamma_{D+1}) = \Gamma_{D+1}, & \epsilon(\Gamma_{D+1}) = 1 \\
\end{array}
\label{10}
\end{equation}

\noindent
For D even the elements $E_{\mu} \;\;\; (\mu=1,\ldots,D)$
and the product $\Gamma_1
\cdots \Gamma_D \Gamma_{D+1}$ are casimirs of CH(D), therefore
in an irreducible representation of CH(D) we get $E_{\mu} =
\eta_{\mu \mu}$, and $\Gamma_{D+1} \sim \Gamma_1 \cdots
\Gamma_D$ which means that the irreps of CH(D) are isomorphic to
those of $C(\eta)$ for all possible signatures of $\eta$ (there
is a unique faithful representation of $C(\eta)$ of dimension $2^D$). For D odd
similar arguments show that the representation theory of CH(D)
is related to that of $C(\eta)$ for $\eta$ a quadratic form
defined in one dimension higher, namely D+1.

The quantum deformation of CH(D), that we will denote
CH$_q$(D), is defined by:
\begin{equation}
\Gamma_{\mu}^2 = \left[ E_{\mu} \right]_q =
\frac{q^{E_{\mu}}-q^{-E_{\mu}}}{q- q^{-1}}
\label{11}
\end{equation}

\noindent
with the rest of equations (\ref{9}) unchanged. The
comultiplication for $\Gamma_{\mu}$ is now given by:
\begin{equation}
\Delta \Gamma_{\mu} = \Gamma_{\mu} \otimes q^{-E_{\mu}/2} +
q^{E_{\mu}/2} \Gamma_{D+1} \otimes \Gamma_{\mu}
\label{12}
\end{equation}

\noindent
The Hopf algebra CH$_q$(D) for $D=2$ is very close to the two
parameter quantum supergroup ${\cal U}_{\alpha, \beta}(su(1,1))$
defined in \cite{R}. The correspondence between both algebras
is given by the substitutions $ q^{E_x} \longrightarrow
\alpha^{E}$ and $ q^{E_y} \longrightarrow
\beta^{E}$. Notice that in $CH_q(2)$ we have two central
elements $E_x, E_y$ and one quantum deformation parameter,
while in ${\cal U}_{\alpha, \beta}(su(1,1))$ there exist one
central element and two parameters. This difference will be
important in the representation theory. We observe that the
``SUSY grading" is played in our case by $\Gamma_3$, and in general
by $\Gamma_{D+1}$ for $D>2$.

Next we proceed to define a sort of affinization of the Hopf
algebra CH$_q$(D). The generators of this new algebra that we
denote $\widehat{CH_q(D)}$ are: $E_{\mu}^{(i)}, \Gamma_{\mu}^{(i)},
\Gamma_{D+1}^{(i)}, \;\;i=0,1$ satisfying (\ref{11}) and (\ref{12})
for each value of $i$.
In what follows we will consider only the case $D=2$\footnote{To define
the algebra $\widehat{CH_q(D)}$ properly we should add to (\ref{11}) the
equivalent to Serre's relations. These will not be relevant for
the discussion in this paper.}.

A two dimensional irrep $\pi_{\xi}$ of $\widehat{CH_q(D)}$
is labelled by three complex parameters $\xi = (z,\lambda_x,\lambda_y) \in
C_{\times}^3$ and reads:
\begin{eqnarray}
\pi_{\xi}(\Gamma^{(0)}_x) =
\left(\frac{\lambda_x^{-1}-\lambda_x}{q-q^{-1}}\right)^{1/2} \left(
\begin{array}{cc} 0 & z^{-1} \\ z & 0 \\ \end{array} \right) &, &
\pi_{\xi}(\Gamma^{(1)}_x) =
\left(\frac{\lambda_x-\lambda_x^{-1}}{q-q^{-1}}\right)^{1/2} \left(
\begin{array}{cc} 0 & z \\ z^{-1} & 0 \\ \end{array} \right)
\nonumber \\
& & \nonumber \\
\pi_{\xi}(\Gamma^{(0)}_y) =
\left(\frac{\lambda_y^{-1}-\lambda_y}{q-q^{-1}}\right)^{1/2} \left(
\begin{array}{cc} 0 & -i z^{-1} \\ i z & 0 \\ \end{array} \right) &, &
\pi_{\xi}(\Gamma^{(1)}_y) =
\left(\frac{\lambda_y-\lambda_y^{-1}}{q-q^{-1}} \right)^{1/2}\left(
\begin{array}{cc} 0 & -i z \\ i z^{-1} & 0 \\ \end{array}
\right) \nonumber \\
& & \nonumber \\
\pi_{\xi}(\Gamma^{(0)}_3) =
\left( \begin{array}{cc} 1 & 0 \\ 0 & -1 \\ \end{array} \right)
&, & \pi_{\xi}(\Gamma^{(1)}_3) =
\left( \begin{array}{cc} 1 & 0 \\ 0 & -1 \\ \end{array} \right)
\label{12'} \\
& & \nonumber \\
\pi_{\xi}(q^{E^{(0)}_x}) = \lambda_x^{-1} & , &
\pi_{\xi}(q^{E^{(1)}_x}) = \lambda_x \nonumber \\
& & \nonumber \\
\pi_{\xi}(q^{E^{(0)}_y}) = \lambda_y^{-1} &, & \pi_{\xi}(q^{E^{(1)}_y}) =
\lambda_y \nonumber
\end{eqnarray}

\noindent
The intertwiner $R_{\xi_1,\xi_2}$ for two of these irreps is
defined by the condition
\begin{equation}
R_{\xi_1 \xi_2} \Delta_{\xi_1 \xi_2}(a) = \Delta_{\xi_2
\xi_1}(a) R_{\xi_1 \xi_2} \;\;\; \forall a \in \widehat{CH_q(2)}
\label{14}
\end{equation}

\noindent
with $\Delta_{\xi_1 \xi_2}= \pi_{\xi_1} \otimes \pi_{\xi_2}
(\Delta)$. Assuming that $R_{\xi_1,\xi_2}$ is an invertible
matrix, then the intertwiner equation (\ref{14}) implies:
\begin{equation}
{\rm tr} \;\Delta_{\xi_1 \xi_2}(a) = {\rm tr} \;\Delta_{\xi_2
\xi_1}(a) \;\;\; \forall a \in \widehat{CH_q(2)}
\label{15}
\end{equation}

\noindent
For $a= \Gamma_x^{(0)} \Gamma_y^{(1)}$ we obtain the following
constraint on the labels of the irreps which admit an intertwiner:
\begin{equation}
\frac{2 (\lambda - \mu)}{(1-\lambda^2)^{1/2}
(1-\mu^2)^{1/2}(z^2 - z^{-2})} = k
\label{16}
\end{equation}

\noindent
with $k$ an arbitrary constant. Equation (\ref{16}) defines a
two dimensional variety embedded in ${\bf C}^{3}$ which can be
uniformized in terms of elliptic functions. Identifying $k$ in
(\ref{16}) with the elliptic modulus we define a new variable
$\varphi$ by:
\begin{equation}
z^2 = cn(\varphi) + i sn(\varphi)
\label{17}
\end{equation}

\noindent
Definig now:
\begin{equation}
\lambda_x = \tanh x \;\; , \;\; \lambda_y = \tanh y
\label{18}
\end{equation}

\noindent
equation (\ref{16}) becomes
\begin{equation}
e^{x-y} = dn(\varphi) + i k sn(\varphi)
\label{19}
\end{equation}

\noindent
which means that $x+y$ is independent of $\varphi$, therefore
each point in the curve (\ref{16}) can be parametrized by two
complex parameters $(\varphi, \psi)$ with $\psi$ defined by
\begin{equation}
\tanh \left( \frac{x+y}{2} \right) = cn(\psi) + i sn(\psi)
\label{20}
\end{equation}

\noindent
The main result of this paper is that, given two irreps lying on
the same curve (\ref{16}),
$\xi_1\;(\varphi_1, \psi_1)$ and $\xi_2\;(\varphi_2, \psi_2)$ the
intertwiner $R$--matrix $R_{\xi_1, \xi_2}$ coincides with the one given in
(\ref{3}) (up to a diagonal change of basis)
provided we identify $u=\varphi_1-\varphi_2$. Notice
from (\ref{17}) that the ``affine" parameter $z$ becomes the
standard exponential in the trigonometric limit. The derivation
of (\ref{3}) is long but straighforward, and we have used the following
identity
among elliptic functions:
\begin{equation}
e(\varphi_1-\varphi_2) = \frac{e(\varphi_1)
(dn(\varphi_1)+1) (dn(\varphi_2)+1) -k^2 e(\varphi_2) sn(\varphi_1)
sn(\varphi_2)}{e(\varphi_2) (dn(\varphi_1)+1) (dn(\varphi_2)+1)
-k^2 e(\varphi_1) sn(\varphi_1) sn(\varphi_2)}
\label{21}
\end{equation}

Summarizing our results we have proved that the intertwiner $R$--matrix for
two dimensional irreps of the Hopf algebra $\widehat{CH_q(2)}$
is the free fermion eight vertex solution to the Yang--Baxter equation.

\section{Comments}

The Sklyanin algebra \cite{S} of the eight vertex model is
determined by the corresponding elliptic curve and the
anisotropy $\gamma$ \cite{Ch}. In the free fermionic case, i.e.
$\gamma = K$, the curve, for the most general case with
non--zero field, is given by (\ref{16}). An important question
that we will address in a future publication, is the
mathematical meaning, inside $\widehat{CH_q(2)}$, of the
$\gamma=K$--Sklyanin algebra.

Taking into account that the trigonometric limit of the free
fermion model is given by the six vertex free fermion model and
that this $R-$matrix is the intertwiner of
$U_{\hat q}(\widehat{gl(1,1)})$, it is plausible to conjecture
that the Hopf algebra $\widehat{CH_q(2)}$ plays the role, in the
sense of reference \cite{FR}, of hidden quantum group of the
$q-$WZW model defined by $U_q(\widehat{gl(1,1)})$. More
precisely we expect that the connection matrices of the $q-$KZ
equation for $U_q(\widehat{gl(1,1)})$ are quantum $6-j$ symbols
of $\widehat{CH_q(2)}$.

Another interesting issue is the interpretation of the eight
vertex free fermion $R-$matrix as an scattering $S-$matrix in
the sense of Zamolodchikov \cite{Z}. From our previous results
we know that the correspondent ``solitons" define now irreducible
representations of $\widehat{CH_q(2)}$. Even though we cannot
expect a field theory limit preserving the elliptic nature of
this $S-$matrix, as a consequence of the c-theorem \cite{ZDV},
the elliptic $S-$matrix (\ref{3}) may still have a good physical
meaning in the lattice, maybe related to the dynamics of the
cnoidal waves in a Toda lattice \cite{T}.

Finally and based on the previously mentioned close connection
between $N\!=\!2$ soliton $S-$matrices and intertwiners of
$U_q(\widehat{gl(1,1)})$, it is natural to wonder if some
relevant information on $N\!=\!2$ integrable models is still
hidden in the quantum Clifford algebra $\widehat{CH_q(2)}$.

\subsection*{Acknowledgements}

We would like A. Berkovich for valuable discussions. The work of
R.C. and E.L.M. is supported by M.E.C. fellowships.

\end{document}